# Multimodal Fusion and Coherence Modeling for Video Topic Segmentation

### v1

@@ Abstract
The video topic segmentation (VTS) task segments videos into intelligible, non-overlapping topics, facilitating efficient comprehension of video content and quick access to specific content. VTS is also critical to various downstream video understanding tasks. Traditional VTS methods using shallow features or unsupervised approaches struggle to accurately discern the nuances of topical transitions. Recently, supervised approaches have achieved superior performance on video action or scene segmentation over unsupervised approaches. In this work, we improve supervised VTS by thoroughly exploring multimodal fusion and multimodal coherence modeling. Specifically, (1) we enhance multimodal fusion by exploring different architectures using cross-attention and mixture of experts. (2) To generally strengthen multimodality alignment and fusion, we pre-train and fine-tune the model with multimodal contrastive learning. (3) We propose a new pre-training task tailored for the VTS task, and a novel fine-tuning task for enhancing multimodal coherence modeling for VTS. We evaluate the proposed approaches on educational videos, in the form of lectures, due to the vital role of topic segmentation of educational videos in boosting learning experiences. Additionally, we introduce a large-scale Chinese lecture video dataset to augment the existing English corpus, promoting further research in VTS. Experiments on both English and Chinese lecture datasets demonstrate that our model achieves superior VTS performance compared to competitive unsupervised and supervised baselines [^1] The code and model checkpoints will be released upon acceptance.
@@ 1 Introduction
Traditional VTS techniques predominantly hinge on shallow features Gandhi et al. (2015); Soares and Barrére (2018b); Ali et al. (2021) and unsupervised methods Gupta et al. (2023) due to the scarcity of labeled data. These methods often fall short in capturing the semantic cues that signal topical shifts in video streams, hence suffer from limited precision. Recent advancements in supervised learning paradigms, benefiting from the availability of large-scale labeled data, have achieved notable performance improvements in various video segmentation tasks, such as video action segmentation Zhou et al. (2018); Tang et al. (2019), scene segmentation Huang et al. (2020); Islam et al. (2023), and topic segmentation Wu et al. (2023); Wang et al. (2023); Xing et al. (2024), surpassing unsupervised methods. Performance can be further enhanced by pre-training on vast volumes of unlabeled data Xu et al. (2021); Mun et al. (2022) or initializing models from pre-trained models Yan et al. (2023) and then fine-tuning the model. Hence, in this work, we focus on further improving supervised methods for VTS.
Compared to text topic segmentation Koshorek et al. (2018); Xing and Carenini (2021); Yu et al. (2023), videos contain rich and diverse multimodal contextual information. Fully utilizing multimodal information, such as visual cues and textual data (e.g., screen text and subtitles), could facilitate more detailed content understanding and in turn more accurate semantic segmentation than relying on text only. Also, coherence is essential for understanding logical structures and semantics. Enhancing coherence modeling has achieved significant improvements in long document topic segmentation Yu et al. (2023). Therefore, we improve supervised VTS methods by thoroughly exploring multimodal fusion and multimodal coherence modeling. We enhance multimodal fusion from the perspectives of model architecture and pre-training and fine-tuning tasks. Specifically, we compare various multimodal fusion architectures built upon Cross-Attention and Mixture-of-Experts (MoE). We investigate the effect of multi-modal contrastive learning for general pre-training and fine-tuning for strengthening cross-modal alignment. For enhancing multimodal coherence modeling, we propose a new pre-training task tailored for the VTS task, and a novel fine-tuning task by elevating intra-topic multimodal feature similarity and inter-topic multimodal feature differences. The proposed approaches are extensively evaluated on educational videos, in the form of lectures, due to the pivotal contributions of topic segmentation of educational videos in bolstering the learning experiences.
@@
  • We propose a supervised multimodal sequence labeling model for VTS, denoted MMVTS model. We compare various multimodal fusion architectures. In addition to applying multimodal contrastive learning for strengthening cross-modal alignment, we propose a new self-supervised pre-training task tailored to the VTS task and a novel fine-tuning task for enhancing multimodal coherence modeling.
@@
Dataset Videos Hours Topics/Video Clips/Topic Tokens/Clip Seconds/Clip Language
NPTEL10 Gandhi et al. (2015) 12 - - - - - English
Videoaula Soares and Barrére (2018a) 44 26.4 - - - - Portuguese
CS80 Soares and Barrére (2019) 80 - - - - - English
MOOC100 Das and Das (2019) 100 100 6.9 - - - English
Coursera37 Chand and Oğul (2021) 37 2.8 16.5 - - - English
AVLecture Gupta et al. (2023) 350 297.5 5.4 46.2 26.6 12.3 English
CLVTS (Ours) 510 395 10.1 35.7 31.2 7.7 Chinese
Table 1: Comparison between our CLVTS and existing course video datasets for video topic segmentation.
@@ 2 Related Work
Text Topic Segmentation The objective of text topic segmentation is to automatically partition text into topically consistent, non-overlapping segments Hearst (1994). By automatically mining clues of topic shifts from large amounts of labeled data Koshorek et al. (2018); Arnold et al. (2019), contemporary supervised models Lukasik et al. (2020); Somasundaran et al. (2020); Zhang et al. (2021); Yu et al. (2023) have demonstrated superior performance compared to unsupervised approaches Riedl and Biemann (2012); Solbiati et al. (2021). Notably, supervised models that excel at modeling long sequences Zhang et al. (2021); Yu et al. (2023) are capable of capturing longer contextual nuances and thereby achieve better topic segmentation performance, compared to models that model local sentence pairs or block pairs Wang et al. (2017); Lukasik et al. (2020). In addition, recent works Somasundaran et al. (2020); Xing et al. (2020); Yu et al. (2023) have demonstrated that strengthening coherence modeling can improve the topic segmentation performance. Therefore, in this work, we explore enhancing coherence modeling for video topic segmentation under the multimodal configurations.
(a) Distribution of Consistency Scores.
(b) Video and Topic Duration (minutes).
(c) Distribution of Subjects.
Figure 1: Statistics of the CLVTS dataset.
Video Topic Segmentation Distinct from text topic segmentation, video topic segmentation is inherently characterized by its multimodal data context; hence, comprehensive utilization of the multimodal information could improve semantic understanding capabilities of a model. Specifically,  Gupta et al. (2023) introduce an unsupervised method that employs the TWFINCH algorithm Sarfraz et al. (2021) to cluster video clips into distinct topics based on learned visual and text features, which are derived from self-supervised tasks of matching the narration with the temporally aligned visual content.  Wang et al. (2023) simply concatenate the sampled encoded visual features and text embedding as the input to a pre-trained language model, which may cause inferior fine-tuning due to the discrepancy between the pre-training stage of the language model and the fine-tuning stage combining textual and visual inputs. Other works related to VTS are summarized and compared in Appendix A.
3 The CLVTS Dataset
Collection Procedure The CLVTS dataset is primarily sourced from Public Video Platforms [^2] https://study.163.com/ [^3] https://www.bilibili.com/v/knowledge. Videos are first transcribed by a competitive automatic speech recognition (ASR) system [^4] https://tingwu.aliyun.com/home. We then ask annotators to combine visual and textual (ASR 1-best) information to mark the timestamp (in seconds) of the end of each topic. We ensure high accuracy and reliability of annotations from three aspects, including annotator training, hierarchical topic labeling, and a multi-annotator strategy. Details of these strategies and the rigorous quality assessment and quality control we conduct can be found in Appendix B.
Dataset Statistics To measure the extent of consensus for VTS annotations, following Shou et al. (2021), we compute the $F_{1}@k$ score (described in Appendix D) based on the absolute distance between two topic boundary sequences, varying the threshold $k$ from 0 to 8 seconds with a step size of 2 seconds, where 8 seconds are approximately the average duration of a clip. By averaging the F1 scores across all three pairs of annotated topic boundaries from three annotators on the same video, we obtain the consistency score. The more similar the annotations from all annotators on the same video are, the higher the consistency score is. Figure 1 shows that the consistency scores of the majority of videos exceed 0.5, indicating a decent degree of consensus for VTS annotations Shou et al. (2021).
Table 1 compares our CLVTS dataset against existing VTS datasets. Among annotated videos in CLVTS, 47% are presentations showing slides, 34% are blackboard demonstrations, and 19% are miscellaneous types. We also collect 1027 hours of unlabeled videos from the same sources for pre-training. Besides the linguistic distinctness, CLVTS is characterized by its natural and uninterrupted long videos, a stark contrast to the English lecture dataset, AVLecture Gupta et al. (2023), since nearly two-thirds of AVLecture are reassembled pre-segmented short videos. Figure 1 and 1 show a diverse distribution of video and topic durations and a broad spectrum of subjects in CLVTS.
@@ 4 Methodology
Figure 2 depicts the architecture of our VTS model. Section 4.1 presents the problem definition of multimodal VTS and the overall architecture. We enhance multimodal fusion from the perspectives of model architecture, pre-training, and fine-tuning tasks. Specifically, we compare different architectures built upon Merge- and Cross-Attention, and Mixture-of-Experts for multimodal fusion (Section 4.1). We explore multimodal contrastive learning for cross-modality alignment and propose a new pre-training task tailored for VTS (Section 4.2). For fine-tuning, we also propose a novel task for multimodal coherence modeling (Section 4.3).
@@
Figure 2: Overall model architecture and four distinct Multimodal Fusion Layer architectures. The visual encoder, including the 2D, 3D and OCR feature extractors, is frozen during our pre-training and fine-tuning. The blue dotted line in $l_{mcssl}$ module denotes the topic boundary. The green solid line in Figure 2 represents the features being brought closer together, while the red dashed line signifies push the features apart.
4.1 Multi-Modal Video Topic Segmentation
Overall Architecture
@@ 4.1 Multi-Modal Video Topic Segmentation
Following prior works Zhang et al. (2021); Wu et al. (2023), we define video topic segmentation as a clip-level sequence labeling task and propose our MultiModal Video Topic Segmentation (MMVTS) model. As illustrated in Figure 2, we apply unimodal pre-trained encoders for each modality and then fuse multimodal information at the intermediate representation level (i.e., ‘middle fusion’ Xu et al. (2023)) through Multimodal Fusion Layers. Given a video, we transcribe it with a competitive automatic speech recognition (ASR) system [^5] https://tingwu.aliyun.com/home and use ASR 1-best as the text modality. We then divide the video into $n$ clips $(c^{v}_{i},c^{t}_{i})_{i=1}^{n}$, with clips segmented at the sentence boundaries predicted on ASR 1-best. $c^{v}_{i}=\{f^{i}_{1},...,f^{i}_{k}\}$ denotes evenly sampled $k$ frames from the $i$-th clip and is fed to a visual encoder $\bm{E_{v}}$ to extract visual features. $c^{t}_{i}=\{w^{i}_{1},...,w^{i}_{\|s_{i}\|+1}\}$ denotes the sequence of words from ASR 1-best within the $i$-th clip, where $w^{i}_{1}$ is the inserted special token [BOS] and $\|s_{i}\|$ denotes the number of words of $i$-th clip. $c^{t}_{i}$ is fed to a text encoder $\bm{E_{t}}$ for extracting textual features. After extracting unimodal features, we first apply trainable projection matrices $W_{v}$ and $W_{t}$ to convert unimodal features into the same dimension, as visual feature sequence $\bm{v}=\{v_{1},...,v_{n}\}$ and textual feature sequence $\bm{t}=\{t_{1},...,t_{n}\}$ (Eq. 1). Then we fuse the multimodal information with $M$ Multimodal Fusion Layers $\bm{MFL}_{M}$ and obtain the updated visual features $\bm{h^{v}}=\{h^{v}_{1},...,h^{v}_{n}\}$ and textual features $\bm{h^{t}}=\{h^{t}_{1},...,h^{t}_{n}\}$ (Eq. 2), which are then concatenated into the multimodal features $\bm{m}=\{m_{1},...,m_{n}\}$ (Eq. 3). Finally, the multimodal features $\bm{m}$ are fed into the predictor composed of a linear layer $W_{p}$ to obtain the probability of binary classification $\bm{p}=\{p_{1},...,p_{n}\}$ (Eq. 4), where $p_{i}$ indicates whether the $i$-th clip is at the boundary of the topic that it belongs to. We use the standard binary cross-entropy loss (Eq. 5) to train the model, where $y_{i}\in\{0,1\}$ is the label. The last clip is excluded from loss computation.
@@
Compared to ‘late fusion’ where no cross-modal interaction happens until after the classifiers, middle fusion and ‘early fusion’ are found to generally outperform late fusion Nagrani et al. (2021), probably because early and middle fusion aligns better with human perception where multimodal fusion happens early in sensory processing. On the other hand, compared to early fusion, middle fusion yields superior or comparable performance Nagrani et al. (2021) and is much less computationally expensive since we could freeze partially strong pre-trained unimodal encoders and only train a small number of parameters in the Multimodal Fusion Layers.
@@
Multimodal Fusion Layer (MFL)
We compare four different cross-modal interaction mechanisms for multimodal fusion layers. We investigate the Merge-Attention and Co-Attention multimodal fusion layers proposed in  Dou et al. (2022); Yang et al. (2023). With Merge-Attention (Figure 2), features from unimodal encoders are sequentially concatenated and then input into a standard transformer encoder layer Vaswani et al. (2017), which shares attention parameters across modalities. A feed forward layer is added on top to produce the final output representation. In contrast, with Co-Attention (Figure 2), features from each unimodal encoder first go through self-attention with modality-specific attention parameters, then we perform symmetrical cross-attention to integrate information from all other modalities to enhance the representation of the considered modality, followed by a feed forward layer.
Inspired by Mustafa et al. (2022), which interleaves MoE encoder layers and standard dense encoder layers for image-text multimodal models, we also investigate two new architectures by replacing the traditional single feed-forward layers in Figure 2 and 2 with a Mixture-of-Experts (MoE) Shazeer et al. (2016); Lepikhin et al. (2020) module respectively, depicted by Figure 2 and 2. The motivation is that adding MoE on top of the fused representations may facilitate deeper cross-modal integration of information and improve model capacity without a proportional increase in computational complexity. Specifically, experts are MLPs activated depending on the input. Firstly, we concatenate fused features output from self-attention or cross-attention. Then, we implement the Noisy Top-k Gating mechanism Shazeer et al. (2016) to select $K$ experts from a total of $E$ candidates (Eq. 68), where $SN()$ denotes stand normal distribution, $W_{n}$ denotes tunable Gaussian noise to help load balancing, $W_{g}$ is a trainable weight matrix, $K$ and $E$ are hyper-parameters. Finally, the outputs of the $K$ activated experts are linearly combined according to the learned gating weights (Eq. 9). For the MoE training objective, we sum importance loss and load loss Shazeer et al. (2016) to balance expert utilization as Eq. 10.
@@
UnsupAVLS Gupta et al. (2023) V+T - 56.00 - 70.86 - - - - - -
@@
PT FT-Coh MMVTS Model Modality
@@
Co-Attn with MoE 59.84 75.62 67.69 72.21 68.84 39.98 58.96 51.41 54.71 51.27
Table 2: Performance of baselines and our MMVTS models on AVLecture and CLVTS test sets. V and T under Modality denote Visual and Textual modality, respectively. MMVTS Baseline1,2,3 denote our MMVTS model w/o Multimodal Fusion Layers. Attn denotes Attention. PT denotes pre-training the model on unlabeled data (Section 4.2 Eq. 13) before fine-tuning. FT-Coh denotes adding the two auxiliary multimodal coherence modeling tasks during fine-tuning (Section 4.3 Eq. 15); w/o FT-Coh refers to fine-tuning with the standard $l_{vts}$ (Eq. 5). For each metric, the best result among all models is bold faced while the best result in each group is underscored.

### v2

@@ Abstract
The video topic segmentation (VTS) task segments videos into intelligible, non-overlapping topics, facilitating efficient comprehension of video content and quick access to specific content. VTS is also critical to various downstream video understanding tasks. Traditional VTS methods using shallow features or unsupervised approaches struggle to accurately discern the nuances of topical transitions. Recently, supervised approaches have achieved superior performance on video action or scene segmentation over unsupervised approaches. In this work, we improve supervised VTS by thoroughly exploring multimodal fusion and multimodal coherence modeling. Specifically, (1) we enhance multimodal fusion by exploring different architectures using Cross-Attention and Mixture of Experts. (2) To generally strengthen multimodality alignment and fusion, we pre-train and fine-tune the model with multimodal contrastive learning. (3) We propose a new pre-training task tailored for the VTS task, and a novel fine-tuning task for enhancing multimodal coherence modeling for VTS. We evaluate our proposed approaches on educational videos, in the form of lectures, due to the vital role of topic segmentation of educational videos in boosting learning experiences. Additionally, to promote research in VTS, we introduce a large-scale Chinese lecture video dataset to augment the existing English lecture video datasets. Experiments on both English and Chinese lecture datasets demonstrate that our model achieves superior VTS performance compared to competitive unsupervised and supervised baselines [^1] The code and model checkpoints will be released upon acceptance.
@@ 1 Introduction
Traditional VTS approaches predominantly hinge on shallow features Gandhi et al. (2015); Soares and Barrére (2018b); Ali et al. (2021) and unsupervised methods Gupta et al. (2023), due to scarcity of labeled data. These methods often fall short in capturing the semantic cues that signal topical shifts in video streams, hence suffer from limited precision. Recent advancements in supervised learning paradigms have achieved notable performance improvements in multi-modal task Yang et al. (2022); Tu et al. (2022, 2023); Zhang et al. (2022) and various video segmentation tasks, such as video action segmentation Zhou et al. (2018); Tang et al. (2019), scene segmentation Huang et al. (2020); Islam et al. (2023), and topic segmentation Wu et al. (2023); Wang et al. (2023); Xing et al. (2024), surpassing unsupervised methods. Performance of supervised approaches can be further enhanced by pre-training on vast volumes of unlabeled data Xu et al. (2021); Mun et al. (2022) or initializing models from pre-trained models Yan et al. (2023) and then fine-tuning the model. Hence, in this work, we focus on further improving supervised methods for VTS.
Compared to text topic segmentation Koshorek et al. (2018); Xing and Carenini (2021); Yu et al. (2023), videos contain rich and diverse multimodal contextual information. Fully utilizing multimodal information, such as visual cues and textual data (e.g., screen text and subtitles), could facilitate more detailed content understanding and in turn more accurate semantic segmentation than relying on text only. Our case studies in Appendix H demonstrate the great challenges posed by VTS, particularly to unsupervised approaches or supervised methods that rely solely on either visual or textual modality. The complexity inherent in video content—where multimodal signals must be effectively integrated—accentuates the difficulty. Also, coherence is essential for understanding logical structures and semantics. Enhancing coherence modeling has achieved significant improvements in long text topic segmentation Yu et al. (2023). Therefore, we improve supervised VTS methods by thoroughly exploring multimodal fusion and multimodal coherence modeling. We enhance multimodal fusion from the perspectives of model architecture and pre-training and fine-tuning tasks. Specifically, we compare various multimodal fusion architectures built upon Cross-Attention and Mixture-of-Experts (MoE). We investigate the effect of multimodal contrastive learning for general pre-training and fine-tuning for strengthening cross-modal alignment. For enhancing multimodal coherence modeling, we propose a new pre-training task tailored for the VTS task, and a novel fine-tuning task by elevating intra-topic multimodal feature similarities and inter-topic multimodal feature differences. The proposed approaches are extensively evaluated on educational videos, in the form of lectures, due to the pivotal contributions of topic segmentation of educational videos in bolstering the learning experiences.
@@
  • We propose a supervised multimodal sequence labeling model for VTS, denoted MMVTS. We explore various multimodal fusion architectures, and apply multimodal contrastive learning for strengthening cross-modal alignment. We also propose a new self-supervised pre-training task tailored to the VTS and a novel fine-tuning task for enhancing multimodal coherence modeling.
@@
Dataset Videos Hours Topics/Video Clips/Topic Seconds/Clip Domain Language Available
NPTEL10 Gandhi et al. (2015) 12 - - - - Education English $\times$
Videoaula Soares and Barrére (2018a) 44 26.4 - - - Education Portuguese
CS80 Soares and Barrére (2019) 80 - - - - Education English
MOOC100 Das and Das (2019) 100 100 6.9 - - Education English $\dagger$
Coursera37 Chand and Oğul (2021) 37 2.8 16.5 - - Education English $\times$
VSTAR Wang et al. (2023) 8159 4625 61.2 0.4 90 Television English $\dagger$
NewsNet Wu et al. (2023) 1000 946 8.5 - - News English $\times$
MultiLive Qiu et al. (2023) 1000 1300 8.8 - - Livestream English $\times$
AVLecture Gupta et al. (2023) 350 297.5 5.4 46.2 12.3 Education English
YouTube Xing et al. (2024) 5422 858.5 6.7 16 5.3 Diverse English $\times$
Behance Xing et al. (2024) 575 1225.2 5.2 248 6.0 Livestream English $\times$
CLVTS (Ours) 510 395 10.1 35.7 7.7 Education Chinese
Table 1: Comparison between our CLVTS dataset and existing video datasets for the video topic segmentation task. $\dagger$ indicates that the data is not entirely open source. Prior to our work, AVLecture is the only publicly available large-scale video dataset supporting supervised VTS methods.
@@ 2 Related Work
Text Topic Segmentation Text topic segmentation aims to automatically partition text into topically consistent, non-overlapping segments Hearst (1994). By automatically mining clues of topic shifts from large amounts of labeled data Koshorek et al. (2018); Arnold et al. (2019), contemporary supervised models Lukasik et al. (2020); Somasundaran et al. (2020); Zhang et al. (2021); Yu et al. (2023) demonstrate superior performance compared to unsupervised approaches Riedl and Biemann (2012); Solbiati et al. (2021). Notably, supervised models that excel at modeling long sequences Zhang et al. (2021); Yu et al. (2023) are capable of capturing longer contextual nuances and thereby achieve better topic segmentation performance, compared to models that model local sentence pairs or block pairs Wang et al. (2017); Lukasik et al. (2020). In addition, recent works Somasundaran et al. (2020); Xing et al. (2020); Yu et al. (2023) show that strengthening coherence modeling can improve text topic segmentation performance. Inspired by these findings, in this work, we explore enhancing coherence modeling for video topic segmentation under the multimodal configurations.
Video Topic Segmentation For video topic segmentation, some approaches, such as BaSSL Mun et al. (2022), explore visual-only information. However, many recent works have achieved enhanced semantic understanding of videos by leveraging multimodal data. Gupta et al. (2023) introduced UnsupAVLS, which uses the TWFINCH algorithm to cluster video clips into topics based on visual and text features. Wang et al. (2023) proposed SWST, which concatenates visual and text features for language models; however, it may suffer from discrepancies between pre-training of the language model and fine-tuning. Wu et al. (2023) focused on hierarchical modeling of scene, story, and topic, without further exploring how to better integrate multimodal features. Xing et al. (2024) employed asymmetric cross-modal attention for obtaining text-aware visual representations. It may be most related to our work. However, our work differs from Xing et al. (2024) as we explore symmetric cross-modal attention and also investigate the Mixture-of-Experts mechanism, as well as introducing topic-level Contrastive Semantic Similarity Learning into fine-tuning for enhanced coherence modeling in the multimodal framework.
@@ 3 Methodology
Figure 1 depicts the overall architecture of our MMVTS model. Section 3.1 presents the problem definition of multimodal VTS and the overall model architecture. We enhance multimodal fusion from the perspectives of model architecture, pre-training, and fine-tuning tasks. Specifically, we compare different multimodal fusion architectures built upon Merge- and Cross-Attention, and Mixture-of-Experts (Section 3.1). We explore multimodal contrastive learning for cross-modality alignment and propose a new pre-training task tailored for VTS (Section 3.2). For fine-tuning, we also propose a novel task for multimodal coherence modeling (Section 3.3).
@@
Figure 1: The overall architecture of our MMVTS model and four distinct architectures of the Multimodal Fusion Layers in (a). In the overall architecture, the snowflake symbol indicates that the parameters of the certain module are frozen; whereas, the flame symbol signifies a trainable module. The blue dotted lines in the $l_{mcssl}$ module denote the topic boundaries. The green solid lines in the $l_{cma}$ module depict the features being brought closer, while the red dashed lines depict the features being pushed apart.
3.1 MultiModal Video Topic Segmentation
Overall Architecture.
@@ 3.1 MultiModal Video Topic Segmentation
Following prior works Zhang et al. (2021); Wu et al. (2023), we define video topic segmentation as a clip-level sequence labeling task and propose our MultiModal Video Topic Segmentation (MMVTS) model. As illustrated in Figure 1, we apply unimodal pre-trained encoders for the vision and text modality, respectively, and then fuse multimodal information at the intermediate representation level (i.e., middle fusion Xu et al. (2023)) through the Multimodal Fusion Layer. Given a video, we transcribe it with a competitive automatic speech recognition (ASR) system [^2] https://tingwu.aliyun.com/home and use ASR 1-best as the text modality. We then divide the video into $n$ clips $(c^{v}_{i},c^{t}_{i})_{i=1}^{n}$, with clips segmented at the sentence boundaries predicted on ASR 1-best. $c^{v}_{i}=\{f^{i}_{1},...,f^{i}_{k}\}$ denotes evenly sampled $k$ frames within the $i$-th clip and is fed into a visual encoder $\bm{E_{v}}$ to extract visual features. $c^{t}_{i}=\{w^{i}_{1},...,w^{i}_{\|s_{i}\|+1}\}$ denotes the sequence of words from ASR 1-best within the $i$-th clip, where $w^{i}_{1}$ is the inserted special token [BOS] and $\|s_{i}\|$ denotes the number of words in the $i$-th clip. $c^{t}_{i}$ is fed into a text encoder $\bm{E_{t}}$ and the last hidden representation of [BOS] for the i-th clip is used as the text representation of the clip. After extracting the unimodal features, we first apply trainable projection matrices $W_{v}$ and $W_{t}$ to convert unimodal features into the same dimension, resulting in the visual feature sequence $\bm{v}=\{v_{1},...,v_{n}\}$ and the textual feature sequence $\bm{t}=\{t_{1},...,t_{n}\}$ (Eq. 1). Then we fuse the multimodal information with $M$ Multimodal Fusion Layers $\bm{MFL}_{M}$ and obtain the updated visual features $\bm{h^{v}}=\{h^{v}_{1},...,h^{v}_{n}\}$ and textual features $\bm{h^{t}}=\{h^{t}_{1},...,h^{t}_{n}\}$ (Eq. 2), which are then concatenated into the multimodal features $\bm{m}=\{m_{1},...,m_{n}\}$ (Eq. 3). Finally, the multimodal features $\bm{m}$ are fed into the predictor consisting of a linear layer $W_{p}$ to obtain the probability of binary classification $\bm{p}=\{p_{1},...,p_{n}\}$ (Eq. 4), where $p_{i}$ indicates whether the $i$-th clip is at a topic boundary. We use the standard binary cross-entropy loss (Eq. 5) to train the model, where $y_{i}\in\{0,1\}$ is the label. The last clip is excluded from loss computation. Considering computational complexity, we freeze the visual encoder while keeping all other parameters trainable.
@@
Compared to late fusion where no cross-modal interaction happens until after independent predictions by each unimodal model, middle fusion and early fusion are found to generally outperform late fusion Nagrani et al. (2021), probably because early and middle fusion aligns better with human perception where multimodal fusion happens early in sensory processing. On the other hand, compared to early fusion, middle fusion yields superior or comparable performance Nagrani et al. (2021) and is much less computationally expensive since we could freeze some strong pre-trained unimodal encoders with a large number of parameters and only train a small number of parameters.
@@
Multimodal Fusion Layer (MFL).
We compare four distinct cross-modal interaction mechanisms for Multimodal Fusion Layers. We investigate the Merge-Attention and Co-Attention multimodal fusion layers proposed in Dou et al. (2022); Yang et al. (2023b). With Merge-Attention (Figure 1), features from unimodal encoders are concatenated sequentially and then input into a standard transformer encoder layer Vaswani et al. (2017), which shares attention parameters across modalities. A feed forward layer is added on top to produce the final output representation. In contrast, with Co-Attention (Figure 1), features from each unimodal encoder first go through self-attention with modality-specific attention parameters, then we perform symmetric cross-attention to integrate information from all other modalities to enhance the representation of the considered modality, followed by a feed forward layer.
Inspired by Mustafa et al. (2022), which interleaves MoE encoder layers and standard dense encoder layers for image-text multimodal models, we also investigate two new architectures by replacing the traditional single feed-forward layers in Figure 1 and 1 with a MoE module Shazeer et al. (2016); Lepikhin et al. (2020). The resulting architectures are depicted by Figure 1 and 1. The motivation is that adding MoE on top of the fused representations may facilitate deeper cross-modal integration of information and improve model capacity without a proportional increase in computational complexity. Specifically, experts are MLPs activated depending on the input. Firstly, we concatenate fused features output from self-attention or cross-attention. Then, we implement the Noisy Top-k Gating mechanism Shazeer et al. (2016) to select $K$ experts from a total of $E$ candidates (Eq. 68), where $SN()$ denotes the standard normal distribution, $W_{n}$ denotes tunable Gaussian noise to help load balancing, $W_{g}$ is a trainable weight matrix, $K$ and $E$ are hyper-parameters. Finally, the outputs of the $K$ activated experts are linearly combined with the learned gating weights (Eq. 9). For the MoE training objective, we sum the importance loss and the load loss Shazeer et al. (2016) to balance expert utilization as in Eq. 10.
@@
UnsupAVLS Gupta et al. (2023) V+T - 56.00$\ddagger$ - 70.86$\ddagger$ - - - - - -
@@
PT FT-Coh MMVTS Models (Ours) Modality
@@
Co-Attn with MoE 59.77 75.01 67.94 71.69 68.61 39.98 58.96 51.41 54.71 51.27
Table 2: Performance of baselines and our MMVTS models on AVLecture and CLVTS test sets. $\ddagger$ denotes the leakage of the ground-truth topic number. V and T under Modality denote Vision and Text modality, respectively. MMVTS Baseline1,2,3 denote our MMVTS model w/o Multimodal Fusion Layers. Attn denotes Attention. PT denotes pre-training the model on unlabeled data (Section 3.2 Eq. 13) before fine-tuning. FT-Coh denotes adding the two auxiliary multimodal coherence modeling tasks during fine-tuning (Section 3.3 Eq. 15); w/o FT-Coh refers to fine-tuning with the standard $l_{vts}$ (Eq. 5). For each metric, the best result among all models is boldfaced while the best result in each group is underscored.